PDS_VERSION = PDS3 RECORD_TYPE = STREAM OBJECT = TEXT PUBLICATION_DATE = 1997-03-07 NOTE = "MGS RST Instrument Health Report HEA7066A.TXT" END_OBJECT = TEXT END Miscellaney =========== Our query for engineering data, with time limits in the PVL file of 1997-03-06 00:00:00-23:59:59 returned 3 points in day 1997-03-07. I believe there's already an ISA in the system on this; but since our ISA viewing capability is also broken, I can't get either the status or number. Our attempt to check epic.jpl.nasa.gov for new EOP files this morning when Netscape was unable to locate the proxy host thorn.jpl.nasa.gov. Perhaps thorn was temporarily down (about 08:10 AM PST). The sendmail process(es) that limited SOPC and NASCOM line throughput earlier this week appear to have been multiplying error messages about undeliverable mail. How they started and why they multiplied are not known. The problem was reported to the SA's yesterday. Engineering Data ================ Engineering Data Statistical Summary from CSV2STAT: First SCET: 1997-03-06T00:07:40.666 Last SCET: 1997-03-07T01:05:36.661 CHANNL CHANNEL TITLE DN-LO DN-HI EU-LO EU-HI EU-AVG/STDEV RECDS ------ ------------- ----- ----- --------- --------- ---------/-------- ----- L-0037 MOT1_USO_ENA 0 0 36449 L-0047 MOT2_USO_ENA 1 1 36449 L-0053 RF_SW_OSC 2 2 36450 L-0060 TWTA1_FILMNT 0 0 36449 L-0061 TWTA1_HV 0 0 36450 L-0070 TWTA2_FILMNT 1 1 36450 L-0071 TWTA2_HV 0 0 36449 L-0080 USO_PWR 0 0 36449 L-0090 KaBLE_STAT 0 1 36449 L-0091 KaBLE_ENABLE 0 0 36450 L-0111 MOT1_RCV_AGC 255 255 -1.52E+02 -1.52E+02 -1.52E+02/0.00E+00 18225 L-0121 MOT2_RCV_AGC 118 255 -1.53E+02 -1.07E+02 -1.14E+02/7.49E-02 18225 L-0131 TWTA1_HLX_I 0 1 -3.84E-02 -1.85E-02 -3.83E-02/8.13E-06 18225 L-0132 TWTA1_ANOD_V 0 2 -1.00E+03 -1.00E+03 -1.00E+03/2.20E-03 18225 L-0141 TWTA2_HLX_I 38 43 7.05E-01 8.05E-01 7.49E-01/1.12E-04 18225 L-0142 TWTA2_ANOD_V 59 64 -8.70E+02 -8.60E+02 -8.66E+02/1.10E-02 18225 L-0200 USO_REG_V 203 205 4.06E+00 4.10E+00 4.08E+00/5.63E-05 18225 L-0201 USO_OVEN_V 123 126 2.46E+00 2.52E+00 2.49E+00/7.43E-05 18225 T-0302 EPC1_T 115 133 2.20E+01 2.98E+01 2.44E+01/2.42E-02 4553 T-0303 EPC2_T 109 138 1.98E+01 3.24E+01 2.26E+01/3.89E-02 4553 T-0310 RF_ISOLTR1_T 179 206 1.73E+01 2.86E+01 2.00E+01/1.67E-02 18225 T-0311 RF_ISOLTR2_T 167 188 2.53E+01 3.28E+01 2.76E+01/1.06E-02 18225 T-0312 TWT1_T 177 207 1.67E+01 2.93E+01 1.96E+01/4.06E-02 4553 T-0313 TWT2_T 160 185 2.64E+01 3.52E+01 2.88E+01/2.54E-02 4553 T-0315 USO_T 174 177 1.09E+01 1.18E+01 1.14E+01/2.55E-03 4553 T-0316 KA_AMP_T 148 199 2.08E+01 3.97E+01 2.44E+01/4.86E-02 9110 L-0200 remains constant at 4.08 volts. L00201 is up 0.01 volts to 2.49 volts, and T-0315 is up 0.6C to 11.4C from yesterday. Open Loop Data ============== Data Collection --------------- No new data Closed Loop Data ================ Data Collection --------------- We have obtained the following files from OSCAR: ATDF: ODF: IONCAL: TROPCAL: WEATHER: UTPM: EOP: Unable to access EPIC (see above)